OpenStudy (anonymous):

How many molecules are in 0.456mol of Al2(SO4)3

OpenStudy (photon336):

This is our molar mass. for aluminum sulfate. \[342.15 \frac{ g }{ mol } \] but there's another rule and it says that 1 mol of any substance has: \[1 mol = (6.02x10^{23}particles)\] set up a proportion, because if we know that this amount of particles is in 1 mol, then we need to find the number of particles in 0.456 moles of our sample. \[\frac{ 1 mol }{ 6.02x10^{23}particles } = \frac{ 0.456mol }{ x particles }\] cross multiply, and we get. \[2.75x10^{23} particles \]

OpenStudy (photon336):

@anabear127